# Online Electron Reconstruction at CLAS12

### v1

@@ 3 Offline Electron Reconstruction in the CLAS12 Forward Detector
The HTCC was built to identify electrons and positrons and separate them from other charged particle types [22]. The HTCC is composed of eight mirrors in each of the six FD sectors. The HTCC mirrors will produce Cerenkov radiation when an electron travels through the mirror, with the radiated photons read out by photo-multiplier tubes (PMTs). Clusters are formed from PMTs collecting light from adjacent mirrors. The number of photo-electrons associated with a particle then becomes a discriminating variable in electron identification. However, pions will start to produce more than two photoelectrons when they have momentum greater than 4.5 to 5 GeV. This effect is clearly demonstrated in Figure 3 which shows an increase above 5 GeV in reconstructed clusters associated to positively charged particles, leading to positive pions being identified as positrons.
@@
The trigger system at CLAS12 uses electron identification in the FD as it main trigger [25]. The requirements placed on electron identification are similar to those in the reconstruction. However, one drawback is that there only is a loose geometrical matching between the DC tracks and hits in other detectors and no calculation of the sampling fraction as the momentum of a track will not be known at the trigger level. The trigger is then susceptible to increased particle misidentification background compared to the reconstruction. The trigger can be run with or without DC roads to improve geometrical matching. The CLAS12 trigger system also allows for other measurements such as photoproduction with low angle electron scattering by requiring an electron in the Forward Tagger and two tracks in the FD. However, the trigger system has some limitations such as not being able to identify two tracks in the same sector. Finally, a J/$\psi$ trigger was added to some CLAS12 run periods that required two minimum ionising particle tracks with low energy deposition in the ECAL and in opposite sectors of the FD.
@@ 4 Online Electron Reconstruction in the CLAS12 Forward Detector
Information from the ECAL cluster can then be passed to a neural network to identify the particle that produced the track and cluster as being an electron or not. The variables used here are the sum of the ADC value in each of the strips, the position of the cluster and the number of strips hit in each view, for each of the PCAL, ECIN and ECOUT. Information from the HTCC is also passed to the electron identification network. As the HTCC only contains 8 PMTs in a sector, the ADC value from all 8 PMTs are passed as input to the neural network. Clusters in the HTCC sometimes leak into neighboring sectors. To account for this, the ADC value from all 8 PMTs in the neighboring sectors are also given to the electron identification network, leading to the HTCC contributing 24 features. The track’s average wire position in each superlayer is also passed to the neural network so that it can correlate the position of HTCC hits to the track. Note that only detector level information such as ADC values are used here to avoid the need for any intermediate reconstruction steps. The training sample for the electron identification algorithm is composed from particles identified as electrons by the offline reconstruction in the positive sample, and all other negative particles in the negative sample. The network has 5 hidden layers with ReLu activation functions and 50, 100, 50, 25 and 5 nodes respectively. The output layer uses a sigmoid activation function.
@@ 5 Validation
Figure 11 shows the online electron identification efficiency and purity relative to the offline electron identification algorithm, within 2 strips of the edge of the calorimeter as shown in Figure 10. The threshold on the response was here placed at 0.075. As shown the electron identification has an efficiency close to 100%. The purity is also above 75% and reaching 90% and above at high momentum. The online electron identification found at least one electron in only 43% of events where the conventional electron trigger with DC roads found at least one electron. Assuming an electron identification efficiency of 100% as shown in Figure 11, the online electron identification could be used as a second pass filter after the conventional electron trigger and achieve a data reduction of 57% in these running conditions. This is comparable to what could be achieved with the CNN based approach of Ref. [1] described in the introduction. It must however be stressed that the method presented here has the added benefit of providing reconstructed electrons for further analysis.
@@
Consideration should also be extended to the photoproduction (so called MesonEx) and J/$\psi$ triggers [25] described at the end of Section 3. The MesonEx trigger requires two charged hadron tracks in the FD. The conventional trigger cannot accommodate this trigger when two tracks are in the same sector. The J/$\psi$ trigger requires two minimum ionising particles (MIPs) in opposite sectors. MIP-candidates are identified as having low energy deposition in the calorimeters. Although no online MIP identification algorithm has been developed, the current scheme can at least reconstruct two tracks in opposite sectors that do not belong to electrons. Figure 12 shows the efficiency in recovering two hadron tracks as a function of the number of sectors separating the two tracks. The efficiency here measures the number of events with at least two tracks with a given sector difference in the online reconstruction for events with at least two tracks with the same sector difference in the offline reconstruction. As shown, the current scheme would allow to accommodate the MesonEx and J/$\psi$ triggers whilst retaining the same efficiency relative to the offline reconstruction algorithm.
Figure 12: The efficiency of the online reconstruction schemes at finding two tracks with a given separation in sectors. The bounding boxes show which sector separation are currently or yet to be implemented in the conventional MesonEx or J/$\psi$ triggers.
Exclusive reactions can be used to identify electrons without having to rely on the offline identification scheme. For example, one can look at the reaction $ep\to e^{\prime}\pi^{+}\pi^{-}(p)$ where the proton is undetected. The events produced in this reaction can be selected using the missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ that will peak at the mass of the neutron (0.94 GeV). The number of $ep\to e^{\prime}\pi^{+}\pi^{-}(p)$ events is counted by fitting the missing mass distribution. The positively charged pion is identified using the offline particle identification scheme. The negatively charged pion candidate is not identified using the offline reconstruction and just required to be associated with a negatively charged track. The electron candidate tracks are required to have at least a hit in one of the calorimeters, and the HTCC is also required to have a hit in the same or neighboring sectors, although the hit does not have to be associated to the electron. The electron identification efficiency can then be obtained by comparing the total number of $ep\to e^{\prime}\pi^{+}\pi^{-}(p)$ events to those where the electron candidate is identified as an electron by either the offline or online reconstruction schemes. Note that the negatively charged pion and the electron tracks are required to be matched in the offline and online reconstructions, ensuring that the efficiency only takes into account the particle identification efficiency.
Figure 13: The missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ for all events selected by the conventional level 1 electron trigger (black) and events where the electron candidate was identified as an electron by the offline (red) and online (blue) identification schemes. The missing mass distributions are plotted with cuts on the distance from the edge of the PCAL at 13.5 cm (left) and without the cuts (right).
@@
Figure 15: Left: The polar angle vs the momentum of negatively charged pions in the missing range of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ from 0.85 to1.0 GeV. Right: The missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ for all events selected by the conventional level 1 electron trigger (black) and events where the pion candidate was not identified as an electron by the offline (red) and online (blue) identification schemes. The missing mass distributions are plotted with cuts on the distance from the edge of the PCAL at 13.5 cm.
@@ 6 Improving the Electron Identification Efficiency
Figure 18: Left: The missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ for all events selected by the conventional level 1 electron trigger (black) and events where the electron candidate was identified as an electron by the offline (red) and online (blue) identification schemes. The missing mass distributions are plotted without cuts on the distance from the edge of the PCAL. For both left and right, the online identification scheme has been re-trained to improve its efficiency. Right: The missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ for all events selected by the conventional level 1 electron trigger (black) and events where the pion candidate was not identified as an electron by the offline (red) and online (blue) identification schemes.

### v2

# A Machine Learning Based Approach to Online Electron Reconstruction at CLAS12

@@ 3 Offline Electron Reconstruction in the CLAS12 Forward Detector
The HTCC was built to identify electrons and positrons and separate them from other charged particle types [22]. The HTCC is composed of eight mirrors in each of the six FD sectors. The HTCC mirrors will produce Cherenkov radiation when an electron travels through the mirror, with the radiated photons read out by photo-multiplier tubes (PMTs). Clusters are formed from PMTs collecting light from adjacent mirrors. The number of photo-electrons associated with a particle then becomes a discriminating variable in electron identification. However, pions will start to produce more than two photoelectrons when they have momentum greater than 4.5 to 5 GeV. This effect is clearly demonstrated in Figure 3 which shows an increase above 5 GeV in reconstructed clusters associated to positively charged particles, leading to positive pions being identified as positrons.
@@
[…]
@@ 4 Online Electron Reconstruction in the CLAS12 Forward Detector
Information from the ECAL cluster can then be passed to a neural network to identify the particle that produced the track and cluster as being an electron or not. The Analog-to-Digital Converter (ADC) value associated to each ECAL hit, representing the amplitude of a continuous signal in an ECAL strip, is used in the offline reconstruction to define the hit energy. The online reconstruction will not benefit from the calibration process that allows to convert between ADC value to energy and so will use the raw ECAL ADC values. For the HTCC, the ADC value is proportional to the number of photoelectrons produced by an electron passing through the HTCC. All ECAL variables used for the online identification are: the sum of the ADC value in each of the strips, the position of the cluster and the number of strips hit in each view, for each of the PCAL, ECIN and ECOUT. Information from the HTCC is also passed to the electron identification network. As the HTCC only contains 8 PMTs in a sector, the ADC value from all 8 PMTs are passed as input to the neural network. Clusters in the HTCC sometimes leak into neighboring sectors. To account for this, the ADC value from all 8 PMTs in the neighboring sectors are also given to the electron identification network, leading to the HTCC contributing 24 features. The track’s average wire position in each superlayer is also passed to the neural network so that it can correlate the position of HTCC hits to the track. Note that only detector level information such as ADC values are used here to avoid the need for any intermediate reconstruction steps. This amounts to a total of 51 input features. The training sample for the electron identification algorithm is composed from particles identified as electrons by the offline reconstruction in the positive sample, and all other negative particles in the negative sample, for a total of 200000 training events and 60000 testing events). The network has 5 hidden layers with ReLu activation functions and 50, 100, 50, 25 and 5 nodes respectively. The output layer uses a sigmoid activation function. The network is implemented using CLAS12 specific Java based tools used in other machine learning applications at CLAS12 [13, 14], available via the Maven build automation tool [26, 27].
@@ 5 Validation
Figure 11 shows the online electron identification efficiency and purity relative to the offline electron identification algorithm, within 2 strips of the edge of the calorimeter as shown in Figure 10. The threshold on the response of the online electron identification algorithm was here placed at 0.075. As shown the electron identification has an efficiency close to 100%. The purity is also above 75% and reaching 90% and above at high momentum. The online electron identification found at least one electron in only 43% of events where the conventional electron trigger with DC roads found at least one electron. Assuming an electron identification efficiency of 100% as shown in Figure 11, the online electron identification could be used as a second pass filter after the conventional electron trigger and achieve a data reduction of 57% in these running conditions. This is comparable to what could be achieved with the CNN based approach of Ref. [1] described in the introduction. It must however be stressed that the method presented here has the added benefit of providing reconstructed electrons for further analysis.
@@
[…]
Exclusive reactions can be used to identify electrons without having to rely on the offline identification scheme. For example, one can look at the reaction $ep\to e^{\prime}\pi^{+}\pi^{-}(p)$ where the proton is undetected and the scattered electron is written as e’. The events produced in this reaction can be selected using the missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ that will peak at the mass of the proton (0.94 GeV). The number of $ep\to e^{\prime}\pi^{+}\pi^{-}(p)$ events is counted by fitting the missing mass distribution. The positively charged pion is identified using the offline particle identification scheme. The negatively charged pion candidate is not identified using the offline reconstruction and just required to be associated with a negatively charged track. The electron candidate tracks are required to have at least a hit in one of the calorimeters, and the HTCC is also required to have a hit in the same or neighboring sectors, although the hit does not have to be associated to the electron. The missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ will only peak at the proton mass if the correct mass hypothesis is assigned to each final state particle. The presence of the proton mass peak ensures that the final state particles were assigned the correct mass, and therefore ensures that the electron and pion candidates are truly an electron and pion respectively. The electron identification efficiency can then be obtained by comparing the total number of $ep\to e^{\prime}\pi^{+}\pi^{-}(p)$ events to those where the electron candidate is identified as an electron by either the offline or online reconstruction schemes. Note that the tracks are required to be matched in the offline and online reconstructions, ensuring that the tracks are found in both the offline and online reconstruction such that the efficiency only takes into account the particle identification efficiency.
Figure 13: The missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ for all events selected by the conventional level 1 electron trigger (black) and events where the electron candidate was identified as an electron by the offline (red) and online (blue) identification schemes. The missing mass distributions are plotted with cuts on the distance from the edge of the PCAL at 13.5 cm (left) and without the cuts (right). The number of electrons estimated from the fit to the proton peak in the missing mass distributions are shown within the legend.
@@
Figure 15: Left: The polar angle vs the momentum of negatively charged pions in the missing range of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ from 0.85 to1.0 GeV. Right: The missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ for all events selected by the conventional level 1 electron trigger (black) and events where the pion candidate was not identified as an electron by the offline (red) and online (blue) identification schemes. The missing mass distributions are plotted with cuts on the distance from the edge of the PCAL at 13.5 cm. The number of electrons estimated from the fit to the proton peak in the missing mass distributions are shown within the legend.
@@ 6 Improving the Electron Identification Efficiency
Figure 18: Left: The missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ for all events selected by the conventional level 1 electron trigger (black) and events where the electron candidate was identified as an electron by the offline (red) and online (blue) identification schemes. Right: The missing mass of $ep\to e^{\prime}\pi^{+}\pi^{-}X$ for all events selected by the conventional level 1 electron trigger (black) and events where the pion candidate was not identified as an electron by the offline (red) and online (blue) identification schemes. For both left and right, the online identification scheme has been re-trained to improve its efficiency. The missing mass distributions are plotted without cuts on the distance from the edge of the PCAL. The number of electrons estimated from the fit to the proton peak in the missing mass distributions are shown within the legend.